# Split Unlearning

### v4

@@ 7 Conclusion
We achieved the first practical Split Unlearning framework with SplitWiper and SplitWiper+, which designed a one-way-one-off propagation to decouple propagation between clients and the server, enabling SISA unlearning with absent clients. SplitWiper+ further enhanced client label privacy against the server by employing DP. Experiments across diverse data distributions and tasks demonstrated that SplitWiper effectively achieved unlearning by involving only the requesting clients, resulting in 0% unlearning accuracy, and 8% better accuracy on retained data than non-SISA unlearning on existing SL frameworks. SplitWiper maintained constant overhead and reduced computational and communication costs by over 99%. SplitWiper+ preserved over 90% of label privacy, effectively protecting sensitive information from the server.
Appendix
Appendix A Proof of Theorem 1
Proof (Theorem 1).
$\forall{V_{a}^{*}},{V_{b}^{*}}\in\mathbb{V}_{exp}$ and $\forall{V}_{t}\in\mathbb{V}$,
$=e^{2\varepsilon}.$
@@ B.1.1 Computational Complexity
  • $\mathcal{H}_{S}$: server-side training complexity; $\mathcal{H}_{out}$: output layer complexity; $\mathcal{I}_{loss}$: size of the server-side loss; $\mathcal{I}_{S}$: size of the server-side outputs; NA: no overhead applicable; $\mathcal{I}_{K}$: size of intermediate outputs; $\mathcal{I}_{Y}$: size of labels ; $\mathcal{W}_{s}$: size of server-side model weights; $\mathcal{D}_{inter}$: size of the client-side intermediate values.
In Algorithm 4, the complexity of the expansion-and-masking scheme $\mathcal{H}_{exp}$ scales linearly with the number of real labels and the size of intermediate values. This complexity remains within the same order of magnitude as $\mathcal{H}_{K}$, especially given the significant size of large models today. This computational load drives clients to use SL in collaboration with a server rather than independently handling the entire learning process. For security and efficiency, the expansion factor $\gamma$ is typically kept small, and training epochs $M$ are much larger, often ranging from several dozen to hundreds. As a result, the computational overhead of the SplitWiper+ framework is substantially lower than that of traditional Vanilla and U-shaped SL frameworks. Moreover, SplitWiper has only $1/M$ of the computational complexity of Vanilla SL, further highlighting its advantages.
Server-side computational complexity. The server’s computational cost for updating its model in SplitWiper and SplitWiper+ is $O(M\cdot\mathcal{H}{S})$, where $M$ denotes the number of training epochs, and $\mathcal{H}{S}$ represents the complexity of a single training iteration. This complexity mirrors that in the Vanilla SL framework, where the server-side complexity is also $O(M\cdot\mathcal{H}_{s})$. Conversely, the U-shaped SL framework generally exhibits slightly lower server-side complexity, which is $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$, because clients undertake output layer operations, thus offloading some of the computational demand from the server.
@@ B.1.3 Server-side Storage Overhead
In SplitWiper, the storage complexity on the server is $O(K\cdot\mathcal{D}{inter}+\mathcal{W}_{s})$, where $K$ is the number of clients and $\mathcal{D}{inter}$ is the average size of the intermediate values transmitted by each client. In SplitWiper+, the intermediate values are expanded, resulting in a storage complexity of $O(K\cdot\gamma\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$ with the average of the label expansion factors $\gamma$.
@@ B.2 Unlearning Scenarios
The communication overhead is determined by transmitting the intermediate outputs from the initiating client $k$ to the server, leading to a complexity of $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ in the proposed frameworks, where $\gamma=1$ represents SplitWiper, and $\gamma>1$ represents SplitWiper+. Additionally, in unlearning scenarios, the server only updates the stored intermediate values related to client $k$ without requiring additional storage costs, allowing us to consider server-side storage overhead during the unlearning scenarios consistent with that of the learning scenarios. It satisfies G1 (independent unlearning), thus improving the robustness and reducing the overhead of unlearning processes.
@@
Appendix C Algorithm of the Frameworks
@@
Appendix D Experiments on Sample-Level Unlearning Tasks
@@
To demonstrate the effectiveness of sample-level unlearning, we focus on the test accuracy of the unlearned label $\mathbb{Y}_{u}$. As shown in Table 7, with increasing $\delta$, the proportion of unlearned samples grows, leading to an accelerated decline in test accuracy for label $\mathbb{Y}_{u}$, while the performance on other labels $\mathbb{Y}_{o}$ remains largely unaffected. These results highlight the ability of our frameworks to effectively meet sample-level unlearning requirements without compromising the utility of other labels.
@@
Appendix E Experiments on Convergence
@@
(Client-side) Training / Transmission Time (s) Accuracy (%)
Existing SL SplitWiper SplitWiper+
Unlearning 0.1   90.58 99.07 98.40 97.15 97.77 96.18 96.22 97.63 97.83
0.5   79.95 90.83 90.37 89.26 89.10 88.63 89.00 89.13 90.08
0.9   63.26 76.40 78.03 71.36 72.37 69.53 70.64 72.53 73.45
  • Notation:Unlearned sample proportion (i.e., $\delta$).
((a)) Perfect clustering accuracy.
((b)) Inclusive clustering accuracy.
Table 12: Training and transmission time consumption (client-side) of different SL frameworks.
Data Distribution Dataset Scenario Client Type Training Time (s) Transmission Time (s)
U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+ U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+
$\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$
Clients with same task Homo labels Hetero volumes CIFAR-10 Learning $C_{u}$$\star$ 74.36 66.03 2.23 9.73 15.85 9.08 216.40 261.63 9.04 18.08 27.13 18.08
$C_{o}$- 58.88 42.65 1.42 6.16 10.25 5.72 172.98 168.02 5.90 11.80 17.70 5.90
Unlearning $C_{u}$$\star$ 34.23 22.18 1.43 5.94 10.20 5.64 167.63 168.72 5.84 11.68 17.52 5.84
$C_{o}$- 57.48 42.89 0.00 0.00 0.00 0.00 172.95 168.42 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 79.88 74.68 1.44 6.31 10.70 5.88 324.58 328.02 8.24 16.48 24.73 24.72
$C_{o}$- 77.21 71.84 1.41 5.98 10.21 5.48 312.58 315.24 6.22 12.44 18.66 6.22
Unlearning $C_{u}$$\star$ 77.03 70.93 1.36 5.92 10.17 5.43 308.95 312.04 6.13 12.25 18.38 6.13
$C_{o}$- 77.15 71.56 0.00 0.00 0.00 0.00 309.45 312.04 0.00 0.00 0.00 0.00
Partially overlapping labels CIFAR-10 Learning $C_{u}$$\star$ 87.58 76.24 2.47 11.07 18.93 9.84 313.31 314.44 10.72 21.44 32.17 21.44
$C_{o}$- 42.99 38.59 1.28 5.53 9.43 5.47 162.61 157.22 5.34 10.68 16.02 10.42
Unlearning $C_{u}$$\star$ 67.88 53.51 1.74 7.72 13.31 6.41 219.54 220.83 7.52 15.04 22.56 15.04
$C_{o}$- 43.49 38.25 0.00 0.00 0.00 0.00 163.04 157.22 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 115.03 113.55 2.26 9.85 17.01 8.62 513.69 516.06 12.72 25.44 38.17 28.17
$C_{o}$- 68.54 60.44 1.08 4.43 7.56 4.19 285.63 287.24 5.28 10.56 15.84 14.32
Unlearning $C_{u}$$\star$ 112.48 110.02 2.16 9.53 16.44 8.04 498.07 500.07 12.40 24.80 37.21 14.40
$C_{o}$- 68.82 62.86 0.00 0.00 0.00 0.00 282.50 284.04 0.00 0.00 0.00 0.00
Hetero labels CIFAR-10 Learning $C_{u}$$\star$ 49.04 43.87 1.44 6.25 10.78 8.87 187.54 189.62 6.40 12.80 19.20 12.80
$C_{o}$- 50.31 44.34 1.46 6.28 10.78 7.37 187.53 189.63 6.40 12.80 19.21 12.82
Unlearning $C_{u}$$\star$ 27.80 22.18 0.73 3.12 5.39 5.57 93.77 96.01 3.20 6.41 9.57 9.60
$C_{o}$- 48.88 43.97 0.00 0.00 0.00 0.00 187.54 189.64 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 79.51 76.99 1.48 6.33 10.79 6.36 312.58 316.04 6.40 12.83 19.22 19.16
$C_{o}$- 79.74 77.02 1.50 6.35 10.83 5.78 312.60 316.03 6.42 12.80 19.22 11.21
Unlearning $C_{u}$$\star$ 78.88 76.66 1.44 6.00 10.26 5.79 296.96 300.04 6.08 12.16 18.24 12.16
$C_{o}$- 79.32 77.11 0.00 0.00 0.00 0.00 312.57 316.04 0.00 0.00 0.00 0.00
Clients with different tasks CIFAR-10 + MNIST Learning $C_{u}$$\star$ 93.61 88.92 2.95 13.50 31.64 13.08 381.11 384.05 12.80 25.61 37.29 24.76
$C_{o}$- 104.97 90.37 3.01 14.55 23.64 13.41 428.04 438.09 14.26 30.13 37.81 28.98
Unlearning $C_{u}$$\star$ 80.97 61.15 2.39 9.89 14.20 9.92 286.96 288.04 9.60 19.20 26.13 19.08
$C_{o}$- 103.79 90.35 0.00 0.00 0.00 0.00 428.09 437.94 0.00 0.00 0.00 0.00
Table 13: Test accuracy of different SL frameworks.
Data Distribution Dataset Scenario Label Type Accuracy (%)
U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+
$\gamma=2$ $\varepsilon=0.1$ $\gamma=2$ $\varepsilon=0.2$ $\gamma=3$ $\varepsilon=0.1$ $\gamma=3$ $\varepsilon=0.2$ $\gamma=\Gamma_{p}$ $\varepsilon=0.1$ $\gamma=\Gamma_{p}$ $\varepsilon=0.2$
Clients with same task Homo labels Hetero volumes CIFAR-10 Learning $Y_{u}$downarrow 95.40 92.70 99.80 98.10 98.70 97.40 97.90 99.40 99.70
$Y_{o}$uparrow 93.47 92.36 97.04 96.67 97.41 97.07 96.16 96.83 97.97
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 94.23 92.94 97.22 96.47 97.91 96.78 97.07 97.14 98.56
CIFAR-100 Learning $Y_{u}$downarrow 91.20 92.30 97.60 97.00 97.10 97.00 98.10 97.40 97.80
$Y_{o}$uparrow 82.01 81.77 85.39 83.75 84.17 82.53 83.44 85.27 86.08
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 82.79 82.02 86.86 83.51 84.95 82.71 84.23 86.19 87.17
Partially overlapping labels CIFAR-10 Learning $Y_{u}$downarrow 95.60 94.30 99.60 99.50 98.40 99.90 99.20 99.60 99.70
$Y_{o}$uparrow 94.03 92.91 97.30 96.34 97.26 95.80 96.94 96.79 97.78
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 95.08 92.35 97.72 97.61 97.19 96.11 96.64 97.81 98.07
CIFAR-100 Learning $Y_{u}$downarrow 91.90 91.50 98.00 97.70 98.00 97.60 97.60 98.00 98.10
$Y_{o}$uparrow 81.98 81.73 83.94 82.47 83.47 82.29 83.42 83.67 84.03
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 83.01 82.11 85.14 83.95 84.51 83.66 84.37 84.86 85.17
Hetero labels CIFAR-10 Learning $Y_{u}$downarrow 93.20 93.90 99.40 99.80 99.80 99.70 99.50 99.90 99.90
$Y_{o}$uparrow 93.04 93.67 98.52 97.15 97.92 95.74 96.92 98.26 98.10
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 93.93 94.13 98.60 96.98 98.53 97.06 97.61 98.59 98.14
CIFAR-100 Learning $Y_{u}$downarrow 90.30 92.10 99.90 99.70 98.90 98.70 99.00 99.00 99.10
$Y_{o}$uparrow 81.42 80.22 87.43 85.52 86.93 84.14 85.77 86.07 87.71
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 82.04 80.95 88.87 86.89 87.93 85.99 87.06 87.59 88.21
Clients with different tasks CIFAR-10 + MNIST Learning $Y_{u}$downarrow 95.80 95.60 99.90 97.20 98.60 96.80 96.80 97.30 97.40
$Y_{o}$uparrow 95.13 95.73 98.64 96.71 97.15 95.09 96.47 96.31 96.71
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 95.12 95.74 99.07 96.89 97.01 95.33 96.34 96.56 96.72

### v6

@@ 7 Conclusion
We achieved the first practical Split Unlearning framework with SplitWiper and SplitWiper+, which designed a one-way-one-off propagation to decouple propagation between clients and the server, enabling SISA unlearning with absent clients. SplitWiper+ further enhanced client label privacy against the server by employing differential privacy.
Experiments across diverse data distributions and tasks demonstrated that SplitWiper effectively achieved unlearning by involving only the requesting clients, resulting in 0% unlearning accuracy, and 8% better accuracy on retained data than non-SISA unlearning on existing SL frameworks. SplitWiper maintained constant overhead and reduced computational and communication costs by over 99%. SplitWiper+ preserved over 90% of label privacy, effectively protecting sensitive information from the server.
@@
Appendix
Appendix A Proof of Theorem 1
Proof (Theorem 1).
$\forall{V_{a}^{*}},{V_{b}^{*}}\in\mathbb{V}_{exp}$ and $\forall{V}_{t}\in\mathbb{V}$,
$=e^{2\varepsilon}.$
@@ B.1.1 Computational Complexity
  • $\mathcal{H}_{S}$: server-side training complexity; $\mathcal{H}_{out}$: output layer complexity; $\mathcal{I}_{loss}$: size of server-side loss; $\mathcal{I}_{S}$: size of server-side outputs; $\mathcal{I}_{K}$: size of intermediate outputs; $\mathcal{I}_{Y}$: size of labels; $\mathcal{W}_{s}$: size of server-side model weights; $\mathcal{D}_{inter}$: size of client-side intermediate values.
In Algorithm 4, the complexity of expansion-and-masking $\mathcal{H}_{exp}$ scales linearly with the number of real labels and the size of intermediate values. This complexity remains within the same order of magnitude as $\mathcal{H}_{K}$, especially given the significant size of large models today. This computational load drives clients to use SL in collaboration with a server rather than independently handling the entire learning process. For security and efficiency, the expansion factor $\gamma$ is typically kept small, and training epochs $M$ are much larger, often ranging from several dozen to hundreds. As a result, the computational overhead of SplitWiper+ is substantially lower than that of traditional Vanilla and U-shaped SL frameworks. Moreover, SplitWiper has only $1/M$ of the computational complexity of Vanilla SL, further highlighting its advantages.
Server-side computational complexity. The server’s computational cost for updating its model in SplitWiper and SplitWiper+ is $O(M\cdot\mathcal{H}{S})$, where $M$ denotes the number of training epochs, and $\mathcal{H}{S}$ represents the complexity of a single training iteration. This complexity mirrors that in the Vanilla SL framework, where the server-side complexity is also $O(M\cdot\mathcal{H}_{s})$. Conversely, the U-shaped SL framework generally exhibits slightly lower server-side complexity, which is $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$, because clients undertake output layer operations, offloading computational demands from the server.
@@ B.1.3 Server-side Storage Overhead
In SplitWiper, server’s storage complexity is $O(K\cdot\mathcal{D}{inter}+\mathcal{W}_{s})$, where $K$ is the number of clients and $\mathcal{D}{inter}$ is the average size of the intermediate values transmitted by each client. In SplitWiper+, the intermediate values are expanded, resulting in the storage complexity of $O(K\cdot\gamma\cdot\mathcal{D}_{inter}+\mathcal{W}_{s})$ with the average of the label expansion factors $\gamma$.
@@ B.2 Unlearning Scenarios
Communication overhead is determined by transmitting the intermediate outputs from the initiating client $k$ to the server, leading to a complexity of $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ in the proposed frameworks, where $\gamma=1$ represents SplitWiper, and $\gamma>1$ represents SplitWiper+. Additionally, in unlearning scenarios, the server only updates the stored intermediate values related to client $k$ without requiring additional storage costs, allowing us to consider server-side storage overhead during unlearning scenarios consistent with that of learning. It satisfies G1 (independent unlearning), improving the robustness and reducing the overhead of unlearning processes.
Appendix C Implementation
@@
Appendix D Evaluating Sample-Level Unlearning Tasks
@@
To demonstrate the effectiveness of sample-level unlearning, we focus on the test accuracy of the unlearned label $\mathbb{Y}_{u}$. With increasing $\delta$ (Table 10), the proportion of unlearned samples grows, leading to an accelerated decline in test accuracy for label $\mathbb{Y}_{u}$, while the performance on other labels $\mathbb{Y}_{o}$ remains largely unaffected. These results highlight the ability to effectively meet sample-level unlearning requirements without compromising the utility of other labels.
@@
(Client-side) Training / Transmission Time (s) Accuracy (%)
Existing SL SplitWiper SplitWiper+
Unlearning 0.1   90.58 99.07 98.40 97.15 97.77 96.18 96.22 97.63 97.83
0.5   79.95 90.83 90.37 89.26 89.10 88.63 89.00 89.13 90.08
0.9   63.26 76.40 78.03 71.36 72.37 69.53 70.64 72.53 73.45
  • Notation:Unlearned sample proportion (i.e., $\delta$).
@@
Appendix E Experiments on Text Modality: AG-News with BERT
To further demonstrate the generality of our framework, SplitWiper and SplitWiper+, we extend our evaluation beyond vision tasks. Unlike methods tailored to specific datasets or architectures, our frameworks are designed to support a wide range of data modalities and can be seamlessly integrated with state-of-the-art models. While Section 6 presents strong performance on standard image benchmarks such as CIFAR-10, CIFAR-100, and MNIST, we supplement our analysis with experiments on a non-vision dataset: AG-News [57], using the mini-BERT [58] architecture.
Following Section 6.1, we conduct class-level unlearning experiments to evaluate the efficiency and effectiveness of our methods. Both SplitWiper and SplitWiper+ (Table 7), achieve 0% accuracy on the unlearned label, indicating complete forgetting. Importantly, during the unlearning phase, only the client requesting unlearning is involved, while all other clients remain unaffected.
((a)) Perfect clustering accuracy.
((b)) Inclusive clustering accuracy.
We further evaluate sample-level unlearning through two lenses: unlearning accuracy and member inference hit rate. As shown in Table 8, both SplitWiper and SplitWiper+ remain effective even when unlearning individual data samples. Compared to traditional SL retraining-based methods, our frameworks achieve competitive unlearning accuracy with reduced cost. Moreover, Table 9 shows that our SplitWiper+ substantially lowers the hit rate of membership inference attacks, approaching the level of U-shaped SL. This highlights the effectiveness of our label protection mechanism in enhancing privacy against inference attacks.
These supplementary experiments validate the generalizability of our frameworks across different data modalities and model architectures. By extending from image to text domains and integrating with different model types like BERT, we demonstrate that SplitWiper and SplitWiper+ are not confined to specific tasks or datasets. The results reinforce practical advantages and flexibility of our design, which makes our frameworks well-suited for real-world applications where model diversity and data heterogeneity are common.
Appendix F Experiments on Label Privacy under BIRCH-based Inference Attacks
To further validate the robustness of SplitWiper+ in protecting label privacy and the privacy of intermediate representations, we extend our evaluation beyond K-Means by introducing BIRCH [59]. BIRCH is a hierarchical clustering method and provides a complementary perspective with different clustering assumptions. This allows us to better assess whether SplitWiper+ can defend against a broader class of unsupervised inference attacks.
To maintain consistency with the clustering-based analysis in Section 6.4, we continue to evaluate label privacy under the BIRCH clustering algorithm using the same two metrics: perfect clustering accuracy and inclusive clustering accuracy. These metrics capture both strict and relaxed privacy leakage scenarios, where perfect clustering requires exact group alignment, and inclusive clustering tolerates partial overlap.
As shown in Figure 8, the results obtained with BIRCH clustering are consistent with those reported using KMeans. Specifically, when no label expansion or differential privacy is applied, such as in SplitWiper, vanilla SL, and U-shaped SL, the attacker can cluster intermediate outputs with near-perfect accuracy. This indicates significant vulnerability in privacy of those frameworks. By contrast, SplitWiper+, which employs label expansion strategies and injects DP noise into intermediate values, substantially reduces both perfect and inclusive clustering accuracy. Notably, perfect clustering accuracy drops to 0% on CIFAR-10 and CIFAR-100, indicating that no original label group could be correctly reconstructed. This result confirms that effectiveness of SplitWiper+ in mitigating clustering-based inference attacks, rendering label quantities and identities unrecognizable.
@@ Appendix F Experiments on Label Privacy under BIRCH-based Inference Attacks
SplitWiper+ is a dedicated framework for protecting the true number of labels, which is an often overlooked yet critical privacy issue in SL. It introduces a novel label expansion strategy that obscures the original label structure, making it significantly harder for the server to infer sensitive class-level information. In combination with DP applied to intermediate values, SplitWiper+ delivers comprehensive privacy protection by simultaneously shielding both semantic label information and representational features. This dual-defense approach positions SplitWiper+ as a robust and generalizable solution for privacy-preserving SL.
Appendix G Evaluating Convergence
@@
Ethics Considerations
Our experiments are conducted on a local testing platform using open-source libraries and public datasets, without connecting to any external or live systems. These experiments do not involve any issues related to animals, human beings, the environment, healthcare, or military factors. As of this writing, these studies have not had real-world impact as they have only been utilized in our experiments. Consequently, we have addressed numerous ethical considerations in our experimental design, strictly adhering to the ethical principles outlined in the Menlo Report.